# Learning to Adapt: Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v1

@@ I-C Comparison Between Robot and Animal Gait Selection
Figure 5: Comparison between animal and robot gait selection policy strategies where across all plots both animal and robot locomote at increasing linear forward velocity. (a) The bottom two plots of all robot data indicates the percentage of each gait utilised at that velocity. Magenta, purple and blue shaded regions indicate the transition phases of $\pi_{G}^{\text{uni}}$, $\pi_{G}^{W_{\text{ext}}}$ and animals respectively. This study compares transition strategies of (top left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\text{CoT}}$ to data collected from dogs [31] and horses [42, 29] in terms of CoT, (top right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ to data collected from horses [33] in terms of foot ground reaction forces, (bottom left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{W_{\text{ext}}}$ to horses [29] in terms of external work, and (bottom right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{c^{\text{err}}}$ to opossums and dogs [31]. (b) Mapping the correlation between different all metrics and the average gait ID selected across all velocities for terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$.
When developing metrics to characterise gait transitions in animals, data is collected over intervals of increasing forward velocity on flat terrain [42, 33, 29, 31]. Hence, within Fig. 5(a) we took the same approach. We also repeat this experiment with $h_{\text{terr}}^{3}$ to investigate correlations between the metrics and the effects of introducing rough terrain, as presented in Fig. 5(b). We also train four additional $\pi_{G}$ policies that individually minimise energy consumption $\pi_{G}^{\text{CoT}}$, actuator-structural forces $\pi_{G}^{\tau_{\%}}$, mechanical work $\pi_{G}^{W_{\text{ext}}}$ and stability $\pi_{G}^{c^{\text{err}}}$, in accordance with Section III-E, to compare their performance with $\pi_{G}^{\text{uni}}$. One unanimous observation across Fig. 5(a) is that animals experience a gait transition phase over a range of velocities [31, 42, 43]. This behaviour is reflected in $\pi_{G}^{\text{uni}}$, where we class a transition phase as where no individual gait occupies more than 75% of the gaits used at a specific speed.
@@ I-C1 Energy Expenditure – Cost of Transport
An animal’s gait transition phase overlaps with the optimal point of transition for CoT, $\lambda^{\text{CoT}}$, to minimise CoT [27, 29], as depicted in Fig. 5(a). This behaviour is reflected by $\pi_{G}^{\text{uni}}$ as it tracks the lowest CoT gait and has a clear transition phase centred around $\lambda^{\text{CoT}}$. However, $\pi_{G}^{\text{CoT}}$ initially fails to establish a transition phase and the transition is earlier than $\lambda^{\text{CoT}}$. This is a result of all $\pi_{G}$ policies experiencing rough terrain during training where a hopping gait has improved CoT efficiency, as presented in Fig. 4. This is further supported by Fig. 5(b), where the introduction of $h_{\text{terr}}^{3}$ leads to a very similar CoT distribution with a much higher variance in gait ID as utility gaits become more effective at metric minimisation. This suggests that training a policy that only considers CoT decreases generality and fails to resemble animal gait selection strategies.
@@ I-C2 Actuator-structural Forces – Foot Contact Forces
Animals are observed to change gait to minimise actuator-structural forces (i.e. musculoskeletal forces) [33], which in biomechanics is measured through ground reaction force, $f_{\text{grf}}$. Similarly, $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ reduce $f_{\text{grf}}$ through selecting the optimal gait for minimising $\tau_{\%}$. This supports that $\tau_{\%}$ is a suitable alternative to $f_{\text{grf}}$, which is further validated by a strong correlation between them within Fig. 5(b). However, not only does $\pi_{G}^{\tau_{\%}}$ maintain a trotting gait past optimal $f_{\text{grf}}$, but also the transition itself is instantaneous. In turn, $\pi_{G}^{\tau_{\%}}$ better reflects the animal data from [33] than $\pi_{G}^{\text{uni}}$. This could be explained by the metric being misinterpreted in [33]; with a high correlation between $\tau_{\%}$ and $f_{\text{grf}}$ with stability metric $c_{\text{avg}}^{\text{err}}$, it suggests instability causes transition, which is rare on flat terrain, as discussed in Section II.
@@ I-C3 Mechanical Work – External Work
Animals are observed to transition to preserve external mechanical work, $W_{\text{ext}}$, but unlike with CoT, the transition happens before $\lambda^{W_{\text{ext}}}$, as shown in Fig. 5(a), demonstrating that minimisation of $W_{\text{ext}}$ is relatively relaxed. $\pi_{G}^{\text{uni}}$ reflects this behaviour as the transition occurs before $\lambda^{W_{\text{ext}}}$ yet minimal $W_{\text{ext}}$ is preserved. While $\pi_{G}^{W_{\text{ext}}}$ can reduce $W_{\text{ext}}$, its transition phase is extended over a larger range of velocities compared to animals, occurring just after $\lambda^{W_{\text{ext}}}$. In turn, this suggests that switching gaits between trotting and running offers minimal reductions in $W_{\text{ext}}$ resulting in a less definitive transition. However, this metric also seems to capture stability due to its high correlation with the stability metrics on $h_{\text{terr}}^{3}$ in Fig. 5(b), providing insight into explaining its reduced role in Fig. 5(a) where only flat terrain is present.
@@ I-C4 Stability – Stride Duration Coefficient of Variation
Within [31] animals are shown to reduce their stride duration coefficient of variation (CV) to preserve stability as high gait periodicity indicates stability. This behaviour is presented in Fig. 5(a), where animals are seen to initiate a transition phase when there is a significant increase in stride CV, consequently leading to a decrease in CV and an increase in stability. $\pi_{G}^{\text{uni}}$ inherits the same strategy as only when a spike in stride CV is experienced does a transition phase begin that results in improved stability through lowering CV. However, this is not the case with $\pi_{G}^{c^{\text{err}}}$ as it acts to reduce stride CV much more aggressively by mixing both slow, fast and auxiliary gaits which results in no clear transition phase being produced.
Overall, only $\pi_{G}^{\text{uni}}$ consistently reflects the behaviour across all animal data sets. In turn, this supports the notion that no singular metric can characterise animal gait selection behaviour and only when unifying them can similar behaviour in robots arise; the minimisation of the metrics is expected and is seen across all $\pi_{G}$ policies, but the intricacies of animal gait transition behaviour are only seen in $\pi_{G}^{\text{uni}}$. Additionally, we have also verified that this behaviour transfers to real-world deployment as explored in Supplementary Information A (Appendix A-A).
Figure 6: Framework deployment in real-world environments to evaluate adaptability (a) The deployment of $\pi_{L}^{\text{bio}}$ on uneven grassy terrain with manual gait selection to cycle through all gaits with a $v_{\text{x}}^{\text{cmd}}$ of $0.5$ m/s where red shaded regions indicate a transition period. For subfigures (b), (c) and (d) our framework is deployed on loose timber, muddy grass and a low-friction board where the event of critical loss of stability is indicated by the red dashed line and the bottom subplot uses the same gait colour code as Fig. 4. (e) Snapshots showing how animals also employ the strategy of using a mixture of auxiliary gaits to overcome challenging terrains, where red circles are swing feet and green circles are stance feet.
@@ I-D Adaption to Real World Terrain
Although we have instilled animal gait transition strategies, gait procedural memory, and adaptive motion into our framework, its real-world proficiency remains uncertain. Grassy terrain may trap swing feet, and the ground often features irregularities. Despite $\pi_{L}^{\text{bio}}$ only observing flat terrain during training, it can deploy all seven gaits on this terrain, as illustrated in Fig. 6(a), demonstrating that gait procedural memory and adaptive motion adjustment successfully transfer to real-world environments, providing a high level of adaptability, as shown in Supplementary Video 4 (Appendix E).
Terrain that causes states of instability presents a significant risk to the robot, hence we test the limits of our framework through deployment on loose timber, muddy grass, and a low-friction board, as presented in Fig. 6(b), Fig. 6(c) and Fig. 6(d) respectively and consolidated in Supplementary Video 5 (Appendix F), with an additional experiment for external perturbations included within Supplementary Information B (Appendix A-B). Each of the presented experiments showcases an off-nominal stability recovery event; however in the nominal scenario, the framework can maintain stability without changing gaits. In the case of loose timber, critical instability is caused when one back foot slips on a plank, causing it to collide with another. In response, $\pi_{G}^{\text{uni}}$ utilises auxiliary gaits pronk and bound to recover, as depicted in Fig. 6(b). This strategy of utilising the auxiliary gaits for stability recovery is seen across all experiments and reflected in animals as highlighted in Fig. 6(e) where a horse is observed to utilise bounding and limping gaits to traverse down complex rock formations.
In all experiments presented, a considerable increase in a combination of $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ is experienced by the robot before a gait transition, while a weaker correlation is observed with CoT; spikes in $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ directly coincide or even preempt gait changes while CoT peaks lag. This is expected with $c_{\text{avg}}^{\text{err}}$ and $W_{\text{ext}}$ as they capture periodicity and base height respectively. However, this is less expected for $\tau_{\%}$ to correlate with stability; this was never a factor investigated within [33], however, the correlation observed in these experiments and in Fig. 5(b) provides strong evidence that this is the case.
@@ II Discussion
From taking inspiration from animal locomotion proficiency attributes, we have developed a locomotion framework capable of traversing complex and high-risk terrain despite the robot not utilising extra-perceptive sensors nor experiencing any rough terrain during the training of $\pi_{L}^{\text{bio}}$. For $\pi_{L}^{\text{bio}}$, this is achieved through including the BGS output $\bm{\beta}_{L}$ (which encodes state dependent pseudo gait procedural memory and adaptive motion adjustments) within the observation space $\bm{o}_{L}$. This proves to be effective at overcoming rough terrain within Fig. 3 as without the presence of $\bm{\beta}_{L}$ within $\bm{o}_{L}$, increased failure and instability are observed. This is the equivalent of removing an animal’s cerebellum functionality, which would result in reduced limb coordination and stability [41], in turn supporting the claim that $\bm{\beta}_{L}$ effectively encodes pseudo gait procedural memory.
$\pi_{G}^{\text{uni}}$ for optimal gait selection greatly expands adaptability through instilling gait selection strategies used by animals. As demonstrated in Fig. 6, $\pi_{G}^{\text{uni}}$ can maintain stability in the event of the terrain undergoing radical structural or friction coefficient adjustments. These scenarios pose risks to robots with vision systems, as they typically cannot detect ground friction or terrain changes beyond their front legs. Through the use of $\pi_{G}^{\text{uni}}$, this limitation is mitigated without implementing resource-heavy extra-perceptive sensors. Comparing Fig. 6(b) through 6(d) to Fig. 6(e) showcases that animals and $\pi_{G}^{\text{uni}}$ utilise multiple auxiliary gaits to prevent failure. This behavior, untargeted during training, suggests that unifying these metrics encodes the intricacies of animal gait transitions.
One provoking observation from this work is that actuator-structural forces appear to characterise instability. Considering that [33] validates by applying increased loads that could cause instability, it suggests this metric was initially misunderstood. Another observation is that despite the employed biomechanics metrics only being tested on animals completing a linear forward trajectory on flat terrain, $\pi_{G}^{\text{uni}}$ upholds animal gait transition strategies across a wide range of terrains and base velocity commands. This supports that these metrics successfully characterise gait transitions and the notion that robots can test biomechanics hypotheses, avoiding the resource, compatibility, and ethical challenges of animal testing. Moving forward, we aim to integrate extra-preceptive sensors for preemptive gait planning to reduce stability risks.
@@ III-A Control Framework Overview
At the core of this work, the Unitree A1 quadruped robot used within all experiments features $12$ degrees of freedom, $n$, which are all modeled as revolute joints, with their angular positions denoted as $\bm{q}\in{\mathbb{R}}^{n}$ and its base orientation represented as a rotation matrix $\bm{R}_{B}\in SO(\text{3})$. As discussed in Section I-A and outlined in Fig. 1, both $\pi_{L}$ and $\pi_{G}$ are integrated within a control framework and supported by the SE and BGS for generation of the robot’s state data and gait references respectively. The final output of $\pi_{L}$ is target joint positions, $\bm{q}^{*}$, which then get converted into joint torques, $\bm{\tau}^{*}$, through the following PD controller to get sent to the motors,

### v2

@@ I-C Comparison Between Robot and Animal Gait Selection
Figure 5: Comparison between animal and robot gait selection policy strategies where across all plots both animal and robot locomote at increasing linear forward velocity. (a) The bottom two plots of all robot data indicates the percentage of each gait utilised at that velocity. Magenta, purple and blue shaded regions indicate the transition phases of $\pi_{G}^{\text{uni}}$, $\pi_{G}^{W_{\text{ext}}}$ and animals respectively. This study compares transition strategies of (top left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\text{CoT}}$ to data collected from dogs [35] and horses [46, 33] in terms of CoT, (top right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ to data collected from horses [37] in terms of foot ground reaction forces, (bottom left) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{W_{\text{ext}}}$ to horses [33] in terms of external work, and (bottom right) $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{c^{\text{err}}}$ to opossums and dogs [35]. (b) Mapping the correlation between different all metrics and the average gait ID selected across all velocities for terrains $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$, with the dataset of each terrain consisting of $31$ data points which themselves are an average of $1000$ samples.
When developing metrics to characterise gait transitions in animals, data is collected over intervals of increasing forward velocity on flat terrain [46, 37, 33, 35]. Hence, within Figure 5(a) we took the same approach. We also repeat this experiment with $h_{\text{terr}}^{3}$ to investigate correlations between the metrics and the effects of introducing rough terrain, as presented in Figure 5(b). We also train four additional $\pi_{G}$ policies that individually minimise energy consumption $\pi_{G}^{\text{CoT}}$, actuator-structural forces $\pi_{G}^{\tau_{\%}}$, mechanical work $\pi_{G}^{W_{\text{ext}}}$ and stability $\pi_{G}^{c^{\text{err}}}$, in accordance with Section III-E, to compare their performance with $\pi_{G}^{\text{uni}}$. One unanimous observation across Figure 5(a) is that animals experience a gait transition phase over a range of velocities [35, 46, 47]. This behaviour is reflected in $\pi_{G}^{\text{uni}}$, where we class a transition phase as where no individual gait occupies more than 75% of the gaits used at a specific speed.
@@ I-C1 Energy Expenditure – Cost of Transport
An animal’s gait transition aligns with the CoT-optimal point, $\lambda^{\text{CoT}}$, to minimise energy cost [5, 33], as shown in Figure 5(a). $\pi_{G}^{\text{uni}}$ mirrors this, tracking the lowest-CoT gait with transitions centred on $\lambda^{\text{CoT}}$. In contrast, $\pi_{G}^{\text{CoT}}$ lacks a defined transition phase and switches earlier, due to training on rough terrain where hopping improves CoT (Figure 4). Figure 5(b) further shows $h_{\text{terr}}^{3}$ induces similar CoT distributions but with greater gait variance, as auxiliary gaits become more effective. This highlights that CoT-only training reduces generality and diverges from natural gait selection.
@@ I-C2 Actuator-structural Forces – Foot Contact Forces
Animals are observed to change gait to minimise actuator-structural forces (i.e. musculoskeletal forces) [37], which in biomechanics is measured through ground reaction force, $f_{\text{grf}}$. Similarly, $\pi_{G}^{\text{uni}}$ and $\pi_{G}^{\tau_{\%}}$ reduce $f_{\text{grf}}$ through selecting the optimal gait for minimising $\tau_{\%}$. This supports that $\tau_{\%}$ is a suitable alternative to $f_{\text{grf}}$, which is further validated by a strong correlation between them within Figure 5(b). However, not only does $\pi_{G}^{\tau_{\%}}$ maintain a trotting gait past optimal $f_{\text{grf}}$, but also the transition itself is instantaneous. In turn, $\pi_{G}^{\tau_{\%}}$ better reflects the animal data from [37] than $\pi_{G}^{\text{uni}}$. This could be explained by the metric being misinterpreted in [37]; with a high correlation between $\tau_{\%}$ and $f_{\text{grf}}$ with stability metric $c_{\text{avg}}^{\text{err}}$, it suggests instability causes transition, which is rare on flat terrain, as discussed in Section II.
@@ I-C3 Mechanical Work – External Work
Animals transition to preserve external mechanical work, $W_{\text{ext}}$, but do so before $\lambda^{W_{\text{ext}}}$, indicating relaxed minimisation (Figure 5(a)). $\pi_{G}^{\text{uni}}$ reflects this behaviour; transition occurs before $\lambda^{W_{\text{ext}}}$ yet minimal $W_{\text{ext}}$ is preserved. While $\pi_{G}^{W_{\text{ext}}}$ can reduce $W_{\text{ext}}$, its transition phase is extended over a larger range of velocities compared to animals, occurring just after $\lambda^{W_{\text{ext}}}$. In turn, this suggests that switching gaits between trotting and running offers minimal reductions in $W_{\text{ext}}$ resulting in a less definitive transition. However, $W_{\text{ext}}$ also seems to capture stability due to its high correlation with the stability metrics on $h_{\text{terr}}^{3}$ in Figure 5(b), providing insight into its reduced role in Figure 5(a) where only flat terrain is present.
@@ I-C4 Stability – Stride Duration Coefficient of Variation
Within [35] animals are shown to reduce their stride duration coefficient of variation (CV) to preserve stability as high gait periodicity indicates stability. This behaviour is presented in Figure 5(a), where animals are seen to initiate a transition phase when there is a significant increase in stride CV, consequently leading to a decrease in CV and an increase in stability. $\pi_{G}^{\text{uni}}$ inherits the same strategy as only when a spike in stride CV is experienced does a transition phase begin that results in improved stability through lowering CV. However, this is not the case with $\pi_{G}^{c^{\text{err}}}$ as it acts to reduce stride CV much more aggressively by mixing both slow, fast and auxiliary gaits which results in no clear transition phase being produced.
Only $\pi_{G}^{\text{uni}}$ consistently reflects all animal data sets and demonstrates the successful instillation of animal gait transition strategies. This also supports the notion that no singular metric can characterise animal gait selection and through unification can similar behaviour in robots arise; the minimisation of the metrics is expected and is seen across all $\pi_{G}$ policies, but the intricacies of animal gait transition strategies are only seen in $\pi_{G}^{\text{uni}}$. Additionally, we have verified this behaviour transfers to real-world deployment within Supplementary Section 2; $\pi_{G}^{\text{uni}}$ achieves an $18\%$ and $30\%$ reduction in CoT compared to trotting and running on grassy terrain while also preventing the failure cases trotting exhibits.
Figure 6: Framework deployment in real-world environments to evaluate adaptability (a) The deployment of $\pi_{L}^{\text{bio}}$ on uneven grassy terrain with manual gait selection to cycle through all gaits with a $v_{\text{x}}^{\text{cmd}}$ of $0.5$ m/s where red shaded regions indicate a transition period. For subfigures (b), (c) and (d) our framework is deployed on loose timber, muddy grass and a low-friction board, with maximum velocity magnitudes of $1$ m/s, $2$ m/s and $1.7$ m/s respectively, where the event of critical loss of stability is indicated by the red dashed line, the bottom subplot uses the same gait colour code as Figure 4, and the heatmaps show the magnitudes of each metric increasing from purple to yellow. (e) Snapshots showing how animals also use a mixture of auxiliary gaits to overcome challenging terrains, where red circles are swing feet and green circles are stance feet.
@@ I-D Adaption to Real World Terrain
Although we have instilled animal gait transition strategies, gait procedural memory, and adaptive motion into our framework, its real-world proficiency remains uncertain. Grassy terrain may trap swing feet, and the ground often features irregularities. Despite $\pi_{L}^{\text{bio}}$ only observing flat terrain during training, it can deploy all seven gaits on this terrain, as illustrated in Figure 6(a), demonstrating that gait procedural memory and adaptive motion adjustment successfully transfer to real-world environments, providing a high level of adaptability, as shown in Supplementary Video S4.
Terrain that causes states of instability presents a significant risk to the robot, hence we test the limits of our framework through deployment on loose timber, muddy grass, and a low-friction board, as presented in Figure 6(b), Figure 6(c) and Figure 6(d) respectively and consolidated in Supplementary Video S5, with an additional experiment for external perturbations included within Supplementary Section 3. Each of the presented experiments showcases an off-nominal stability recovery event; however in the nominal scenario, the framework can maintain stability without changing gaits. In the case of loose timber, critical instability is caused when one back foot slips on a plank, causing it to collide with another. In response, $\pi_{G}^{\text{uni}}$ utilises auxiliary gaits pronk and bound to recover, as depicted in Figure 6(b). This strategy of utilising the auxiliary gaits for stability recovery is seen across all experiments and reflected in animals as highlighted in Figure 6(e) where a horse is observed to utilise bounding and limping gaits to traverse down complex rock formations.
In all experiments presented, a considerable increase in a combination of $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ is experienced before a gait transition, while a weaker correlation is observed with CoT; spikes in $W_{\text{ext}}$, $\tau_{\%}$ and $c_{\text{avg}}^{\text{err}}$ directly coincide or even preempt gait changes while CoT peaks lag. This is expected with $c_{\text{avg}}^{\text{err}}$ and $W_{\text{ext}}$ as they capture periodicity and base height respectively. However, this is less expected for $\tau_{\%}$ to correlate with stability; this was never a factor investigated within [37], however, the correlation observed in these experiments and in Figure 5(b) provides strong evidence that this is the case.
@@ II Discussion
From taking inspiration from animal locomotion proficiency attributes, we have developed a locomotion framework capable of traversing complex and high-risk terrain despite the robot not utilising exteroceptive sensors nor experiencing any rough terrain during the training of $\pi_{L}^{\text{bio}}$. For $\pi_{L}^{\text{bio}}$, this is achieved through including the BGS output $\bm{\beta}_{L}$ (which encodes state dependent pseudo gait procedural memory and adaptive motion adjustments) within the observation space $\bm{o}_{L}$. This proves to be effective at overcoming rough terrain within Figure 3 as without the presence of $\bm{\beta}_{L}$ within $\bm{o}_{L}$, increased failure and instability are observed. This is the equivalent of removing an animal’s cerebellum functionality, which would result in reduced limb coordination and stability [45], in turn supporting the claim that $\bm{\beta}_{L}$ effectively encodes pseudo gait procedural memory.
$\pi_{G}^{\text{uni}}$ for optimal gait selection greatly expands adaptability through instilling gait selection strategies used by animals. As demonstrated in Figure 6, $\pi_{G}^{\text{uni}}$ can maintain stability in the event of the terrain undergoing radical structural or friction coefficient adjustments. These scenarios pose risks to robots with vision systems, as they typically cannot detect ground friction or terrain changes beyond their front legs. Through the use of $\pi_{G}^{\text{uni}}$, this limitation is mitigated without implementing resource-heavy exteroceptive sensors. Comparing Figure 6(b) through 6(d) to Figure 6(e) showcases that animals and $\pi_{G}^{\text{uni}}$ utilise multiple auxiliary gaits to prevent failure. This behavior, untargeted during training, suggests that unifying these metrics encodes the intricacies of animal gait transitions.
One provoking observation is that actuator-structural forces appear to characterise instability. Considering that [37] validates by applying increased loads that could cause instability, it suggests this metric was initially misunderstood. Another observation is that despite the employed biomechanics metrics only being tested on animals completing a linear forward trajectory on flat terrain, $\pi_{G}^{\text{uni}}$ upholds animal gait transition strategies across a wide range of terrains and base velocity commands. This supports that these metrics successfully characterise gait transitions and the notion that robots can test biomechanics hypotheses, avoiding the resource, compatibility, and ethical challenges of animal testing. Moving forward, we aim to integrate our work with those that focus on Froude-free locomotion, such as [11, 12, 29] to achieve exemplary adaptability and efficiency at both Froude-characterised and Froude-free locomotion levels.
@@ III-A Control Framework Overview
At the core of this work, the Unitree A1 quadruped robot used within all experiments features $12$ degrees of freedom, $n$, which are all modeled as revolute joints, with their angular positions denoted as $\bm{q}\in{\mathbb{R}}^{n}$ and its base orientation represented as a rotation matrix $\bm{R}_{B}\in SO(\text{3})$. As discussed in Section I-A and outlined in Figure 1, both $\pi_{L}$ and $\pi_{G}$ are integrated within a control framework and supported by the SE and BGS for generation of the robot’s state data and gait references respectively. The final output of $\pi_{L}$ is target joint positions, $\bm{q}^{*}$, which then get converted into joint torques, $\bm{\tau}^{*}$, through the following PD controller to get sent to the motors,